Get the procedure to the answers shown below  Motor M rotates with an angular velocity of w = a(1 - ebt)Determine the magnitudes of the velocity and acceleration of point
Answer #1

Given Motor M rotates with angular velocity wn=all-eo Given a = 40ls - b = -1/ W = 4(1-et) rls radius=s, radius=82 lot let wslan 1-6 angular velocity of fan w angulan acceleration dt angular acceleration of fan - 01 vclocity of point P, V = 8 L 8 = dI

acceleration ap = 10.0149 iris 10.02 in)s velocity (at any time) of point P, Up = 8 3 Wp = 16 (1-6) inls - At To find man velI hope you understand. Comment if you have any doubt.